# Movable Antenna Enhanced Networked Full-Duplex Integrated Sensing and Communication System

### v2

@@ II-A System Model
And, the channel between the $k$-th DL user and the $l$-th UL user can be expressed by
${h}_{du,k,l}(\mathbf{r}_{0,k},\mathbf{r}_{1,l})^{*}=(\mathbf{h}^{3}_{k,l}(\mathbf{r}_{0,k}))^{H}\mathbf{\Sigma}^{2}_{k,l}\mathbf{h}^{2}_{k,l}(\mathbf{r}_{1,l}).$ (11)
@@
Besides the steering vector of the $m$-th TBS towards the sensing target and/or clutters is represented by
@@
where index 00 represent the sensing target, $\mathbf{a}^{4}_{m,j}\triangleq[\cos(\theta^{t,4}_{m,j})\sin(\phi^{t,4}_{m,j});\sin(\theta^{t,4}_{m,j})]^{T}$, and $\theta^{t,4}_{m,j}$ and $\phi^{t,4}_{m,j}$ denote the elevation and azimuth AoDs between TBS $m$ and target/clutters. Similarly, the steering vector of the $p$-th RBS towards the sensing target/clutters also can be written as
@@
where $\mathbf{a}^{5}_{p,j}\triangleq[\cos(\theta^{r,4}_{p,j})\sin(\phi^{r,4}_{p,j});\sin(\theta^{r,4}_{p,j})]^{T}$, and $\theta^{r,4}_{p,j}$ and $\phi^{r,4}_{p,j}$ denote the elevation and azimuth AoAs between RBS $p$ and target/clutters. Therefore, the $m$-th TBS-target/clutters and the $p$-th RBS-target/clutters channels are respectively given as
@@
where $\{\beta^{t}_{m,j}\}$ and $\{\beta^{r}_{p,j}\}$ are complex fading coefficients and are assumed to be known.
Denote $\mathbf{s}\triangleq[s_{1},\cdots,s_{k},\cdots,s_{K_{d}}]^{T}\in\mathbb{C}^{K_{d}\times 1}$ and $\mathbf{s}_{r}\in\mathbb{C}^{N_{t}\times 1}$ as the communication symbol transmitted to DL users and radar probing signal from TBS, respectively. For simplicity, it is assumed that $\mathbb{E}\{\mathbf{s}\mathbf{s}^{H}\}=\mathbf{I}$, $\mathbb{E}\{\mathbf{s}_{r}\mathbf{s}_{r}^{H}\}=\mathbf{I}$ and $\mathbb{E}\{\mathbf{s}\mathbf{s}_{r}^{H}\}=\mathbf{0}$. The transmit signal of the $m$-th TBS can be written as
@@
where $s_{l}^{u}$ and $q_{l}$ denote the information symbol and transmission power of $l$-th UL user, respectively, and $s_{l}^{u}$ are mutually uncorrelated and each has zero mean and unit variance.
The received signal at the DL user $k$ can be written as
$y_{D,k}=\underbrace{{\sum}_{m=1}^{M_{t}}\mathbf{h}_{d,m,k}^{H}\mathbf{x}_{m}}\limits_{\textrm{DL signals}}+\underbrace{{\sum}_{l=1}^{K_{u}}{h}_{du,k,l}x^{u}_{l}}\limits_{\textrm{UL signals}}+n_{d,k}$ (17)
@@
where $\hat{\mathbf{h}}_{d,k}\triangleq[\mathbf{h}_{d,1,k}^{T},\mathbf{h}_{d,2,k}^{T},\cdots,\mathbf{h}_{d,M_{t},k}^{T}]^{T}\in\mathbb{C}^{M_{t}N_{t}\times 1}$, $\hat{\mathbf{w}}_{k}\triangleq[\mathbf{w}_{1,k}^{T},\mathbf{w}_{2,k}^{T},\cdots,\mathbf{w}_{M_{t},k}^{T}]^{T}\in\mathbb{C}^{M_{t}N_{t}\times 1}$, $\bar{\mathbf{W}}_{r}\triangleq[(\mathbf{W}^{r}_{1})^{T},(\mathbf{W}^{r}_{2})^{T},\cdots,(\mathbf{W}^{r}_{M_{t}})^{T}]^{T}\in\mathbb{C}^{M_{t}N_{t}\times N_{t}}$, $\hat{\mathbf{W}}_{r}\triangleq\mathrm{blkdiag}(\bar{\mathbf{W}}_{r})$ and $n_{d,k}\sim\mathcal{CN}(0,\sigma_{d,k}^{2})$ is denoted as the complex additive white Gaussian noise (AWGN) at the $k$-th DL user.
The received signal at $p$-th RBS can be expressed as
@@
$+\underbrace{{\sum}_{l=1}^{K_{u}}\mathbf{h}_{u,p,l}x^{u}_{l}}\limits_{\textrm{UL signals}}+\mathbf{n}_{r,p},$
where $\mathbf{n}_{r,p}\sim\mathcal{CN}(0,\sigma_{r}^{2}\mathbf{I})$ is the AWGN at the $p$-th RBS, $\alpha_{t,j}$ represents the target radar cross section (RCS) between TBS and RBS and $\alpha_{t,j}\sim\mathcal{CN}(0,\sigma_{t,j}^{2})$.
To obtain the information of UL users and target separately, RBS $p$ will adopt $K_{u}+1$ linear filters $\mathbf{u}_{p,i}\in\mathbb{C}^{M_{r}\times 1}$, $\forall i\in\mathcal{J}_{u}\triangleq\{0\}\cup\mathcal{K}_{u}$, to post-process the received signal, where $\mathbf{u}_{p,0}$ corresponds to the radar sensing filter. Therefore, the output information of $i$-th filter at $p$-th RBS can be written as
@@
The achievable rate of $k$-th DL user and $l$-th UL user can be obtained as
@@
$\frac{|\hat{\mathbf{h}}_{d,k}^{H}\hat{\mathbf{w}}_{k}|^{2}}{{\sum}_{i\neq k}^{K_{d}}|\hat{\mathbf{h}}_{d,k}^{H}\hat{\mathbf{w}}_{i}|^{2}+\|\hat{\mathbf{h}}_{d,k}^{H}\hat{\mathbf{W}}^{r}\|^{2}_{F}+{\sum}_{l=1}^{K_{u}}q_{l}|{h}_{du,k,l}|^{2}+\sigma_{d,k}^{2}},$
and the SINR for each UL user can be written as (25), as shown on the top of this page.
$\textrm{SINR}_{u,l}=\frac{q_{l}|\hat{\mathbf{u}}_{l}^{H}\hat{\mathbf{h}}_{u,l}|^{2}}{\sum_{i\neq l}^{K_{u}}q_{i}|\hat{\mathbf{u}}_{l}^{H}\hat{\mathbf{h}}_{u,i}|^{2}+\sum_{j=0}^{K_{t}}\sum_{k=1}^{K_{d}}\sigma_{t,j}^{2}|\hat{\mathbf{u}}_{l}^{H}\hat{\mathbf{g}}_{r,j}\hat{\mathbf{g}}_{t,j}^{H}\hat{\mathbf{w}}_{k}|^{2}+\sum_{j=0}^{K_{t}}\sigma_{t,j}^{2}\|\hat{\mathbf{u}}_{l}^{H}\hat{\mathbf{g}}_{r,j}\hat{\mathbf{g}}_{t,j}^{H}\hat{\mathbf{W}}_{r}\|^{2}_{F}+\sigma_{r}^{2}\|\hat{\mathbf{u}}_{l}^{H}\|^{2}_{2}}$ (25)
The radar SINR [22] for sensing can be derived as (26), as shown on the top of next page.
$\textrm{SINR}_{t}=\frac{\sum_{k=1}^{K_{d}}\sigma_{t,0}^{2}|\hat{\mathbf{u}}_{0}^{H}\hat{\mathbf{g}}_{r,0}\hat{\mathbf{g}}_{t,0}^{H}\hat{\mathbf{w}}_{k}|^{2}+\sigma_{t,0}^{2}\|\hat{\mathbf{u}}_{0}^{H}\hat{\mathbf{g}}_{r,0}\hat{\mathbf{g}}_{t,0}^{H}\hat{\mathbf{W}}_{r}\|^{2}_{F}}{\sum_{i=1}^{K_{u}}q_{i}|\hat{\mathbf{u}}_{0}^{H}\hat{\mathbf{h}}_{u,i}|^{2}+\sum_{j=1}^{K_{t}}\sum_{k=1}^{K_{d}}\sigma_{t,j}^{2}|\hat{\mathbf{u}}_{0}^{H}\hat{\mathbf{g}}_{r,j}\hat{\mathbf{g}}_{t,j}^{H}\hat{\mathbf{w}}_{k}|^{2}+\sum_{j=1}^{K_{t}}\sigma_{t,j}^{2}\|\hat{\mathbf{u}}_{0}^{H}\hat{\mathbf{g}}_{r,j}\hat{\mathbf{g}}_{t,j}^{H}\hat{\mathbf{W}}_{r}\|^{2}_{F}+\sigma_{r}^{2}\|\hat{\mathbf{u}}_{0}^{H}\|^{2}_{2}}$ (26)
II-B Problem Formulation
Our goal is to maximize the sum-rate of all DL and UL users via jointly optimizing the transmit beamformer $\{\mathbf{w}_{m,k},\mathbf{W}^{r}_{m}\}$, the linear filters $\{\mathbf{u}_{p,i}\}$, the DL users’ uplink transmit power $\{q_{l}\}$, the positions of all MAs $\{\mathbf{t}^{0}_{m,n},\mathbf{t}^{1}_{p,j},\mathbf{r}_{0,k},\mathbf{r}_{1,l}\}$. The sum-rate optimization problem can be formulated as
[…]
@@ II-B Problem Formulation
where $\Gamma_{r}$ denotes the minimum target sensing SINR threshold, $P^{BS}_{m}$ and $P_{u,l}$ represent the maximum transmission power of the $m$-th TBS and the $l$-th UL user, respectively, and $D_{t}$ is the minimum distance between MAs of BSs to avoid the coupling effect. Obviously, due to the non-convex objective function and constraints, the optimization problem (P0) is highly challenging.
@@ III-A Problem Reformulation
To make the problem (P0) more tractable, we first employ the weighted minimum mean squared error (WMMSE) method [23] to equivalently transform the objective functions $\textrm{R}_{d,k}$ and $\textrm{R}_{u,l}$ into (29) and (30) , respectively, at the top of next page, where $\{\omega_{d,k}\}$, $\{\beta_{d,k}\}$, $\{\omega_{u,l}\}$ and $\{\beta_{u,l}\}$ are the introduced auxiliary variables. Therefore, the original problem (P0) can be reexpressed as
[…]
@@
$\tilde{\mathrm{R}}_{d,k}\triangleq\textrm{log}(\omega_{d,k})-\omega_{d,k}\big{(}1-2\textrm{Re}\{\beta_{d,k}^{\ast}\hat{\mathbf{h}}_{d,k}^{H}\hat{\mathbf{w}}_{k}\}+|\beta_{d,k}|^{2}({\sum}_{i=1}^{K_{d}}|\hat{\mathbf{h}}_{d,k}^{H}\hat{\mathbf{w}}_{i}|^{2}+\|\hat{\mathbf{h}}_{d,k}^{H}\hat{\mathbf{W}}^{r}\|^{2}_{F}+{\sum}_{l=1}^{K_{u}}q_{l}|{h}_{du,k,l}|^{2}+\sigma_{d,k}^{2})\big{)}+1$ (29)
@@
$+|\beta_{u,l}|^{2}\big{(}{\sum}_{i=1}^{K_{u}}q_{i}|\hat{\mathbf{u}}_{l}^{H}\hat{\mathbf{h}}_{u,i}|^{2}+{\sum}_{j=0}^{K_{t}}\sigma_{t,j}^{2}({\sum}_{k=1}^{K_{d}}|\hat{\mathbf{u}}_{l}^{H}\hat{\mathbf{g}}_{r,j}\hat{\mathbf{g}}_{t,j}^{H}\hat{\mathbf{w}}_{k}|^{2}+\|\hat{\mathbf{u}}_{l}^{H}\hat{\mathbf{g}}_{r,j}\hat{\mathbf{g}}_{t,j}^{H}\hat{\mathbf{W}}_{r}\|^{2}_{F})+\sigma_{r}^{2}\|\hat{\mathbf{u}}_{l}^{H}\|^{2}_{2}\big{)}\bigg{)}+1$

### v3

@@ II-A Channel Model
The FRVs of the transmit and receive channel paths between DL user $k$ and UL user $l$ are respectively expressed as
$\mathbf{h}^{2}_{k,l}(\mathbf{r}_{1,l})=[e^{j\frac{2\pi}{\lambda}\rho^{t,2}_{k,l,1}},e^{j\frac{2\pi}{\lambda}\rho^{t,2}_{k,l,2}},\cdots,e^{j\frac{2\pi}{\lambda}\rho^{t,2}_{k,l,L^{t,2}_{k,l}}}]^{T},$ (11)
$\mathbf{h}^{3}_{k,l}(\mathbf{r}_{0,k})=[e^{j\frac{2\pi}{\lambda}\rho^{r,2}_{k,l,1}},e^{j\frac{2\pi}{\lambda}\rho^{r,2}_{k,l,2}},\cdots,e^{j\frac{2\pi}{\lambda}\rho^{r,2}_{k,l,L^{r,2}_{k,l}}}]^{T},$ (12)
where $\rho^{t,2}_{k,l,i}(\mathbf{r}_{1,l})\triangleq\mathbf{r}_{1,l}^{T}\mathbf{a}^{t,2}_{k,l,i}=x^{u}_{l}\cos(\theta^{t,2}_{k,l,i})\sin(\phi^{t,2}_{k,l,i})+y^{u}_{l}\sin(\theta^{t,2}_{k,l,i})$ and $\rho^{r,2}_{k,l,i}(\mathbf{r}_{0,k})\triangleq\mathbf{r}_{0,k}^{T}\mathbf{a}^{r,2}_{k,l,i}=y^{d}_{k}\sin(\theta^{r,2}_{k,l,i})+x^{d}_{k}\cos(\theta^{r,2}_{k,l,i})\sin(\phi^{r,2}_{k,l,i})$. And, the channel between the $k$-th DL user and the $l$-th UL user can be written as
${h}_{du,k,l}^{*}(\mathbf{r}_{0,k},\mathbf{r}_{1,l})=(\mathbf{h}^{3}_{k,l}(\mathbf{r}_{0,k}))^{H}\mathbf{\Sigma}^{2}_{k,l}\mathbf{h}^{2}_{k,l}(\mathbf{r}_{1,l}),$ (13)
where the matrix $\mathbf{\Sigma}^{2}_{k,l}$ is denoted as the response of all transmit and receive paths from the $l$-th UL user to the $k$-th DL user.
@@
b) Radar Channel: The steering vector of the $m$-th TBS towards the sensing target and/or clutters is represented by
@@
where index 0 represents the sensing target, $\mathbf{a}^{4}_{m,j}\triangleq[\cos(\theta^{t,4}_{m,j})\sin(\phi^{t,4}_{m,j});\sin(\theta^{t,4}_{m,j})]^{T}$, and $\theta^{t,4}_{m,j}$ and $\phi^{t,4}_{m,j}$ denote the elevation and azimuth AoDs between TBS $m$ and target/clutters, respectively. Similarly, the steering vector of RBS $p$ towards the sensing target/clutters can also be written as
@@
where $\mathbf{a}^{5}_{p,j}\triangleq[\cos(\theta^{r,4}_{p,j})\sin(\phi^{r,4}_{p,j});\sin(\theta^{r,4}_{p,j})]^{T}$, and $\theta^{r,4}_{p,j}$ and $\phi^{r,4}_{p,j}$ denote respectively the elevation and azimuth AoAs between RBS $p$ and target/clutters. Therefore, the $m$-th TBS-target/clutters and the $p$-th RBS-target/clutters channels are respectively given as
@@
where $\{\beta^{t}_{m,j}\}$ and $\{\beta^{r}_{p,j}\}$ are complex fading coefficients and are assumed to be known [1].
II-B Signal Model
Denote $\mathbf{s}\triangleq[s_{1},\cdots,s_{k},\cdots,s_{K_{d}}]^{T}\in\mathbb{C}^{K_{d}\times 1}$ and $\mathbf{s}_{r}\in\mathbb{C}^{N_{t}\times 1}$ as the communication symbol transmitted to DL users and radar probing signal from TBS, respectively. For simplicity, it is assumed that $\mathbb{E}\{\mathbf{s}\mathbf{s}^{H}\}=\mathbf{I}$, $\mathbb{E}\{\mathbf{s}_{r}\mathbf{s}_{r}^{H}\}=\mathbf{I}$ and $\mathbb{E}\{\mathbf{s}\mathbf{s}_{r}^{H}\}=\mathbf{0}$. The transmit signal of the $m$-th TBS can be written as  [^3] The joint communication and sensing signal model effectively exploits the available DoF [37].
@@ II-B Signal Model
where $s_{l}^{u}$ and $q_{l}$ denote the information symbol and transmission power of UL user $l$, respectively, and $s_{l}^{u}$ are mutually uncorrelated and each has zero mean and unit variance.
The received signal at DL user $k$ can be written as
$y_{D,k}=\underbrace{{\sum}_{m=1}^{M_{t}}\mathbf{h}_{d,m,k}^{H}\mathbf{x}_{m}}\limits_{\textrm{DL users signals}}+\underbrace{{\sum}_{l=1}^{K_{u}}{h}_{du,k,l}x^{u}_{l}}\limits_{\textrm{UL users signals}}+n_{d,k}$ (19)
@@
where $\hat{\mathbf{h}}_{d,k}\triangleq[\mathbf{h}_{d,1,k}^{T},\mathbf{h}_{d,2,k}^{T},\cdots,\mathbf{h}_{d,M_{t},k}^{T}]^{T}\in\mathbb{C}^{M_{t}N_{t}\times 1}$, $\hat{\mathbf{w}}_{k}\triangleq[\mathbf{w}_{1,k}^{T},\mathbf{w}_{2,k}^{T},\cdots,\mathbf{w}_{M_{t},k}^{T}]^{T}\in\mathbb{C}^{M_{t}N_{t}\times 1}$, $\bar{\mathbf{W}}_{r}\triangleq[(\mathbf{W}^{r}_{1})^{T},(\mathbf{W}^{r}_{2})^{T},\cdots,(\mathbf{W}^{r}_{M_{t}})^{T}]^{T}\in\mathbb{C}^{M_{t}N_{t}\times N_{t}}$, $\hat{\mathbf{W}}_{r}\triangleq\mathrm{blkdiag}(\bar{\mathbf{W}}_{r})$ and $n_{d,k}\!\!\sim\!\!\mathcal{C\!N}(0,\sigma_{d,k}^{2})$ is denoted as the complex additive white Gaussian noise (AWGN) at DL user $k$.
II-C Problem Formulation
a) Communication Rate: The received signal at the $p$-th RBS can be expressed as
@@ II-C Problem Formulation
$+\underbrace{{\sum}_{l=1}^{K_{u}}\mathbf{h}_{u,p,l}x^{u}_{l}}\limits_{\textrm{UL users signals}}+\mathbf{n}_{r,p},$
where $\mathbf{n}_{r,p}\sim\mathcal{CN}(0,\sigma_{r}^{2}\mathbf{I})$ is the AWGN at the $p$-th RBS, $\alpha_{t,j}$ represents the target RCS between TBS and RBS and $\alpha_{t,j}\sim\mathcal{CN}(0,\sigma_{t,j}^{2})$ [38].
To obtain the information of UL users and target separately, RBS $p$ will adopt $K_{u}+1$ linear filters $\mathbf{u}_{p,i}\in\mathbb{C}^{M_{r}\times 1}$, $\forall i\in\mathcal{J}_{u}\triangleq\{0\}\cup\mathcal{K}_{u}$, to post-process the received signal, where $\mathbf{u}_{p,0}$ corresponds to the radar sensing filter. Therefore, the output information of the $i$-th filter at the $p$-th RBS can be written as
@@
The achievable rate of the $k$-th DL user and the $l$-th UL user can be obtained as
@@
$\frac{|\hat{\mathbf{h}}_{d,k}^{H}\hat{\mathbf{w}}_{k}|^{2}}{{\sum}_{i\neq k}^{K_{d}}|\hat{\mathbf{h}}_{d,k}^{H}\hat{\mathbf{w}}_{i}|^{2}+\|\hat{\mathbf{h}}_{d,k}^{H}\hat{\mathbf{W}}^{r}\|^{2}_{2}+{\sum}_{l=1}^{K_{u}}q_{l}|{h}_{du,k,l}|^{2}+\sigma_{d,k}^{2}},$
and the SINR formulation of UL user is written in (27), as shown on the top of this page.
$\textrm{SINR}_{u,l}=\frac{q_{l}|\hat{\mathbf{u}}_{l}^{H}\hat{\mathbf{h}}_{u,l}|^{2}}{\sum_{i\neq l}^{K_{u}}q_{i}|\hat{\mathbf{u}}_{l}^{H}\hat{\mathbf{h}}_{u,i}|^{2}+\sum_{j=0}^{K_{t}}\sum_{k=1}^{K_{d}}\sigma_{t,j}^{2}|\hat{\mathbf{u}}_{l}^{H}\hat{\mathbf{g}}_{r,j}\hat{\mathbf{g}}_{t,j}^{H}\hat{\mathbf{w}}_{k}|^{2}+\sum_{j=0}^{K_{t}}\sigma_{t,j}^{2}\|\hat{\mathbf{u}}_{l}^{H}\hat{\mathbf{g}}_{r,j}\hat{\mathbf{g}}_{t,j}^{H}\hat{\mathbf{W}}_{r}\|^{2}_{2}+\sigma_{r}^{2}\|\hat{\mathbf{u}}_{l}^{H}\|^{2}_{2}}$ (27)
b) Sensing SINR: In the target detection problem, since the detection probability for a given false alarm probability monotonically increases with the sensing SINR, we select the sensing SINR as the metric [39]-[40]. The radar SINR [39]-[40] for sensing can be derived as (28), as shown on the top of this page.
$\textrm{SINR}_{t}=\frac{\sum_{k=1}^{K_{d}}\sigma_{t,0}^{2}|\hat{\mathbf{u}}_{0}^{H}\hat{\mathbf{g}}_{r,0}\hat{\mathbf{g}}_{t,0}^{H}\hat{\mathbf{w}}_{k}|^{2}+\sigma_{t,0}^{2}\|\hat{\mathbf{u}}_{0}^{H}\hat{\mathbf{g}}_{r,0}\hat{\mathbf{g}}_{t,0}^{H}\hat{\mathbf{W}}_{r}\|^{2}_{F}}{\sum_{i=1}^{K_{u}}q_{i}|\hat{\mathbf{u}}_{0}^{H}\hat{\mathbf{h}}_{u,i}|^{2}+\sum_{j=1}^{K_{t}}\sum_{k=1}^{K_{d}}\sigma_{t,j}^{2}|\hat{\mathbf{u}}_{0}^{H}\hat{\mathbf{g}}_{r,j}\hat{\mathbf{g}}_{t,j}^{H}\hat{\mathbf{w}}_{k}|^{2}+\sum_{j=1}^{K_{t}}\sigma_{t,j}^{2}\|\hat{\mathbf{u}}_{0}^{H}\hat{\mathbf{g}}_{r,j}\hat{\mathbf{g}}_{t,j}^{H}\hat{\mathbf{W}}_{r}\|^{2}_{2}+\sigma_{r}^{2}\|\hat{\mathbf{u}}_{0}^{H}\|^{2}_{2}}$ (28)
c) Objective Function and Constraints: Our goal is to maximize the sum-rate of all DL and UL users via jointly optimizing the transmit beamformer $\{\mathbf{w}_{m,k},\mathbf{W}^{r}_{m}\}$, the linear filters $\{\mathbf{u}_{p,i}\}$, the DL users’ uplink transmit power $\{q_{l}\}$, the positions of all MAs $\{\mathbf{t}^{0}_{m,n},\mathbf{t}^{1}_{p,j},\mathbf{r}_{0,k},\mathbf{r}_{1,l}\}$. The weighted sum-rate optimization problem can be formulated as
$\textrm{(P0)}:$ $\mathop{\textrm{max}}\limits_{\{\mathbf{w}_{m,k},\mathbf{W}^{r}_{m}\},\{\mathbf{u}_{p,i}\},\{q_{l}\},\atop\{\mathbf{t}^{0}_{m,n},\mathbf{t}^{1}_{p,j},\mathbf{r}_{0,k},\mathbf{r}_{1,l}\}}\ \!\!\!{\sum}_{k=1}^{K_{d}}\mu_{D,k}\textrm{R}_{d,k}\!\!+\!\!{\sum}_{l=1}^{K_{u}}\mu_{U,l}\textrm{R}_{u,l}$ (29a)
@@
where $\mu_{D,k}\in(0,1)$ and $\mu_{U,l}\in(0,1)$, and ${\sum}_{k=1}^{K_{d}}\mu_{D,k}+{\sum}_{l=1}^{K_{u}}\mu_{U,l}=1$  [^4] Since the weight factors are treated as constants in the optimization process, we will omit them in the following., $\Gamma_{r}$ denotes the minimum target sensing SINR threshold, $P^{BS}_{m}$ and $P_{u,l}$ represent the maximum transmission power of the $m$-th TBS and the $l$-th UL user, respectively, and $D_{t}$ is the minimum distance between MAs of BSs to avoid the coupling effect. Obviously, due to the non-convex objective function and constraints, the optimization problem (P0) is highly challenging.
@@ III-A Problem Reformulation
To make the problem (P0) more tractable, we first employ the WMMSE method [34] to equivalently transform the objective functions $\textrm{R}_{d,k}$ and $\textrm{R}_{u,l}$ into (31) and (32), respectively, at the top of the next page, where $\{\omega_{d,k}\}$, $\{\beta_{d,k}\}$, $\{\omega_{u,l}\}$ and $\{\beta_{u,l}\}$ are the introduced auxiliary variables. Therefore, the original problem (P0) can be reexpressed as
$\textrm{(P1)}:$ $\mathop{\textrm{max}}\limits_{\{\mathbf{w}_{m,k},\mathbf{W}^{r}_{m}\},\{\mathbf{u}_{p,i}\},\{q_{l}\},\atop\{\mathbf{t}^{0}_{m,n},\mathbf{t}^{1}_{p,j},\mathbf{r}_{0,k},\mathbf{r}_{1,l}\}}\ \!\!{\sum}_{k=1}^{K_{d}}\tilde{\mathrm{R}}_{d,k}\!+\!{\sum}_{l=1}^{K_{u}}\tilde{\mathrm{R}}_{u,l}$ (30a)
@@
$\tilde{\mathrm{R}}_{d,k}\triangleq\textrm{log}(\omega_{d,k})-\omega_{d,k}\big(1-2\textrm{Re}\{\beta_{d,k}^{\ast}\hat{\mathbf{h}}_{d,k}^{H}\hat{\mathbf{w}}_{k}\}+|\beta_{d,k}|^{2}({\sum}_{i=1}^{K_{d}}|\hat{\mathbf{h}}_{d,k}^{H}\hat{\mathbf{w}}_{i}|^{2}+\|\hat{\mathbf{h}}_{d,k}^{H}\hat{\mathbf{W}}^{r}\|^{2}_{2}+{\sum}_{l=1}^{K_{u}}q_{l}|{h}_{du,k,l}|^{2}+\sigma_{d,k}^{2})\big)+1$ (31)
@@
$+|\beta_{u,l}|^{2}\big({\sum}_{i=1}^{K_{u}}q_{i}|\hat{\mathbf{u}}_{l}^{H}\hat{\mathbf{h}}_{u,i}|^{2}+{\sum}_{j=0}^{K_{t}}\sigma_{t,j}^{2}({\sum}_{k=1}^{K_{d}}|\hat{\mathbf{u}}_{l}^{H}\hat{\mathbf{g}}_{r,j}\hat{\mathbf{g}}_{t,j}^{H}\hat{\mathbf{w}}_{k}|^{2}+\|\hat{\mathbf{u}}_{l}^{H}\hat{\mathbf{g}}_{r,j}\hat{\mathbf{g}}_{t,j}^{H}\hat{\mathbf{W}}_{r}\|^{2}_{2})+\sigma_{r}^{2}\|\hat{\mathbf{u}}_{l}^{H}\|^{2}_{2}\big)\bigg)+1$